OpenStudy (anonymous):

a)If z= a+b i find z^2 in the form X+ Y i b)hence show that (z^2)* = (z*) ^2 for all complex numbers z c) repeat a and b but for z^3 instead of z^2

OpenStudy (anonymous):



OpenStudy (jamesj):

So what do you have for part a)?

OpenStudy (jamesj):

\[ z^2 = (a + bi)^2 = a^2 + 2abi + b^2i^2 = ... what? \]

OpenStudy (anonymous):

^He's right. Just distribute terms and recognize that since \(i \equiv \sqrt{-1}\) then it follows that \(i^2=-1\). For part b, recall the definition of a complex conjugate: its' a reflection about the real axis. Thus, just change every \(i\) you see to a \(-i\). Generically, the number \(c+di\) for constants \(c\) and \(d\) will have the complex conjugate \(c-di\). (Complex conjugates are denoted with a star (like you put) or a bar above). Simplify both algebraically as you would normally to prove. If you can get that far, try part c, and if you need help, just holler.